OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. –18 B. –12 C. –30 D. –24

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (godlovesme):

plug 3 for x and try to solve :) @AidanWithanAnota

OpenStudy (anonymous):

D?

OpenStudy (godlovesme):

how did u get that ?

OpenStudy (anonymous):

7*3 + 2*3 - 3 ) 7*3=21 + 6 = 27 - 3 = 24

OpenStudy (godlovesme):

\[3-7(3)+2(3)\] right?

OpenStudy (anonymous):

yep

OpenStudy (godlovesme):

\[3-12+6 = -18+6= -12\] \[got \it?\]

OpenStudy (godlovesme):

got it**

OpenStudy (anonymous):

Oh B? and could you help me with 2 more questions?

OpenStudy (godlovesme):

sure :)